OpenStudy (anonymous):

If a crowd of people decided to make a human pyramid, with one person on the top row, two on the second, three on the third, etc, and there are 149 rows, how many people participated?

OpenStudy (anonymous):

arithmetic sequence...

OpenStudy (anonymous):

Okay, so you got yourself here a 2 step problem. first off you need to find the last term. So, you can do taht with one simple formula, and that is a(n)=(n-1)d, where d is 1 and n is 149

OpenStudy (anonymous):

So, solve for all that, and you get 148=a(n) (read a sub n), soooo, now you have another formula... : S=n/2 (a1+a(n)). Now, substitute 149/2 (148+1), and you geeeet 149